OpenStudy (anonymous):

simplify sq root 3 (cube root 3)

OpenStudy (trantom):

1.44224957031 is the cube root of 3... ? is that what you want?

OpenStudy (campbell_st):

rewrite them in index form \[3^{\frac{1}{2}} \times 3^{\frac{1}{3}}\] the rule for multiplying the same base is add the powers

OpenStudy (jdoe0001):

as campbell_st said keep in mind that \(\bf \Large { \sqrt[m]{a^n}=a^{\frac{n}{m}}\qquad thus\\ \quad \\ \sqrt{3}\cdot \sqrt[3]{3}\implies \sqrt[2]{3}\cdot \sqrt[3]{3}\implies 3^{\frac{1}{2}}\cdot 3^{\frac{1}{3}}}\)

OpenStudy (anonymous):

can't figure this out~ options are a. 3 6/5 b. 243 d. 3 5/6. I already tried c. 3 1/5

OpenStudy (jdoe0001):

well, why not add the fractions? recall that exponent rules when you have the same base, you add their exponents and keep the base

OpenStudy (anonymous):

lol